OpenStudy (anonymous):

1 + cot^2(-θ)? sin^2(-θ) + cos^2(-θ) ?

OpenStudy (anonymous):

Do you know that \[\sin^2(x) + \cos^2(x) = 1\] ?

OpenStudy (anonymous):

yes, but sin^2(-θ) + cos^2(-θ) = -sin^2 + cos^2 so how do i figure that out?